the formula for the perimeter of a rectangle is p=2l+2w where l is the length and w is the width. a rectangle has a perimeter of 24 inches. find its dimensions if its length is three inches greater than its width.
Answer by stanbon(75887)

the formula for the perimeter of a rectangle is p=2l+2w where l is the length and w is the width.
a rectangle has a perimeter of 24 inches.
find its dimensions if its length is three inches greater than its width
----
Let width = w
The length = x+3
----------------------
P = 2(L+W)
24 = 2(w+3+w)
12 = 2w+3
2w = 9
w = 4.5 inches (width)
---
length = 4.5+3 = 7.5 inches (length)
======================================
Cheers,
Stan H.